# Adversarial Distilled Retrieval-Augmented Guarding Model for Online Malicious Intent Detection

### v1

@@ 1 Introduction
Beyond these limitations, most traditional methods are designed for offline use and cannot effectively support real-time, context-aware intent detection due to architectural and latency constraints. For example, a music application query such as “songs like ‘I Deserve to Bleed’”, intended for search and recommendation, may be incorrectly flagged as unsafe if the model lacks knowledge that “I Deserve to Bleed” is a song title. In practice, enabling such capabilities requires overcoming two main challenges: (1) lightweight models may meet latency requirements but often lack the reasoning and generalization needed for nuanced or adversarial queries; and (2) effective moderation increasingly requires dynamic, input-specific context that static classifiers cannot handle.
In this paper, we present ADRAG (Adversarial Distilled Retrieval-Augmented Guard), a unified framework for accurate, robust, and efficient online malicious intent detection. During training, a high-capacity teacher model is trained on adversarially perturbed, retrieval-augmented inputs to learn context-aware classification under adversarial conditions. During inference, a distillation encoder scheduler transfers the teacher’s knowledge to a compact student guard model, which is continually updated with an online knowledge base. At deployment, the student model retrieves top-$K$ safety examples from the knowledge base to enable real-time, context-aware malicious detection.
@@
    We demonstrate that ADRAG, with a 149M-parameter model, matches the performance of GPT-4, WildGuard-7B, and Llama-Guard-3-8B across ten safety benchmarks, while sustaining an online latency under 9 ms at 300 QPS.
@@ 2 Related Work
Retrieval-augmented generation (RAG) enhances grounding and robustness by incorporating external context (Lewis et al., 2020). Beyond generation, recent work has extended RAG for safety tasks. Class-RAG (Chen et al., 2024) fine-tune Llama-3-8b with retrieved examples to adapt to changing moderation policies, while GuardAgent (Xiang et al., 2024) leverages retrieval to enforce safety constraints in planning agents. However, retrieved content may introduce harmful or misleading information (An et al., 2025). While RAG improves adaptability, its retrieval component introduces new latency challenges. To address this, Knowledge distillation (Hinton et al., 2015) trains compact student models to approximate the behavior of larger teacher models. This approach has enabled efficient NLP models such as DistilBERT (Sanh et al., 2019), and MiniLM (Wang et al., 2020). In the AI safety domain, distillation has been applied to build lightweight intent detection systems, such as HarmAug (Zhang et al., 2024c). Our work combines RAG’s adaptability with distillation’s efficiency to deliver accurate, robust, and controllable malicious intent detection for real-time deployment.
@@ 3.1 Large Language Models
$F_{\theta}(\mathbf{y}\mid\mathbf{x})=\prod_{t=n+1}^{T}p(x_{t}\mid x_{<t},\mathbf{x}),$ (1)
@@ 3.2 Retrieval-Augmented Generation
$\mathcal{D}_{\mathbf{x}}=\mathcal{R}_{K}(\mathbf{x};\mathcal{D}),$ (2)
@@ 3.3 Adversarial Training
$\min_{\theta}\;\mathbb{E}_{\mathbf{x},\mathbf{y}}\left[\max_{\|\delta\|\leq\rho}-\log F_{\theta}^{\delta}(\mathbf{y}\mid\mathbf{x})\right],$ (3)
@@ 3.4 Knowledge Distillation
$\mathcal{L}_{\text{distill}}=\mathrm{KL}\left(F_{\theta_{T}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\,\|\,F_{\theta_{S}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\right),$ (4)
@@ 3.5 Problem Statement
$\hat{y}=F_{\theta}(\mathbf{x},\mathcal{D}_{\mathbf{x}}),\quad\mathcal{T}_{\text{tot}}(\mathbf{x})=\mathcal{T}_{\text{ret}}(\mathbf{x})+\mathcal{T}_{\text{inf}}(\mathbf{x},\mathcal{D}_{\mathbf{x}})\leq\tau,$ (5)
where $\mathcal{T}_{\text{tot}}$ denotes the total system latency, comprising retrieval latency $\mathcal{T}_{\text{ret}}$ and inference latency $\mathcal{T}_{\text{inf}}$, which need to remain below a strict threshold $\tau$ to enable real-time moderation. To support context-aware prediction, the query $\mathbf{x}$ is concatenated with $K$ retrieved examples: $\mathbf{x};\mathbf{d}_{\mathbf{x}}^{(1)};\ldots;\mathbf{d}_{\mathbf{x}}^{(K)}]$, where each $\mathbf{d}_{\mathbf{x}}^{(i)}=(\bar{\mathbf{x}}_{i},\bar{y}_{i})$ represents a prior query and its associated intent label. The concatenated input is fed into the classifier $F_{\theta}$ (or $\mathbf{x}$ alone if $\mathcal{D}_{\mathbf{x}}=\emptyset$), with prompt engineering is used to ensure a coherent sequence that preserves semantic clarity and guides model behavior. This context-enhanced formulation improves reliability by leveraging semantically similar examples, while maintaining the strict latency constraints required for real-time applications.
@@ 4.1 Overview
We propose ADRAG, a two-stage framework for Accurate, Robust, and Efficient online malicious intent detection, as illustrated in Figure 1. In the training phase, a high-capacity teacher model $F_{\theta_{T}}(\mathbf{x},\mathcal{D}_{\mathbf{x}})$ is optimized via Retrieval-Augmented Adversarial Fine-Tuning (RAFT), where each input $\mathbf{x}$ is paired with a retrieved context set $\mathcal{D}_{\mathbf{x}}$ that injects external, often domain-specific knowledge to enrich under-specified queries, substantially enhancing the model’s predictive accuracy, especially for ambiguous or information-sparse cases. To further enhance robustness, ADRAG simulates real-world noise in the knowledge base, retrieval pipeline, context quality, and user input through adversarial perturbations, incorporating them into training via an adversarial loss $\mathcal{L}_{\text{adv}}$ that promotes consistent and reliable predictions under degraded conditions. For deployment, ADRAG applies Selective Knowledge Distillation (SKD) to compress the teacher into a lightweight student model $F_{\theta_{S}}$, guided by an encoder scheduler that governs training dynamics by assigning each epoch to operate in teacher-only, co-distillation, or student-focused mode. At inference time, the student retrieves top-$K$ relevant contexts $\mathcal{D}_{\mathbf{x}}\subset\mathcal{D}$ for each query $\mathbf{x}$ to enable accurate, low-latency predictions that generalize beyond the static training distribution and remain resilient to adversarial or ambiguous inputs, while adaptability is maintained through an Evolving Knowledge Base (EKB) continuously updated with user feedback and policy-driven synthetic examples.
@@ 4.2.1 Supervised Contextual Fine-Tuning
$\mathcal{D}_{\mathbf{x}_{i}}=\left\{\mathbf{d}\in\mathcal{R}_{K}(\mathbf{x}_{i};\mathcal{D})\;\middle|\;\text{sim}(f_{\text{enc}}(\mathbf{x}_{i}),f_{\text{enc}}(\mathbf{d}))\geq 1-\epsilon\right\},$ (6)
@@
$\mathcal{L}_{\text{train}}=\frac{1}{N}\sum_{i=1}^{N}\text{CE}\left(F_{\theta_{T}}(\mathbf{x}_{i},\mathcal{D}_{\mathbf{x}_{i}}),y_{i}\right),$ (7)
@@
$\mathcal{D}_{\mathbf{x}_{i}}^{\Delta}=\mathcal{R}_{K}(\mathbf{x}_{i};\mathcal{D},f_{\text{enc}}^{\Delta},\text{sim}^{\Delta}),$ (8)
@@
$\mathcal{D}_{\mathbf{x}_{i}}^{\Delta}=\left\{\mathbf{d}\in\mathcal{D}\;\middle|\;\delta\leq\text{sim}(f_{\text{enc}}(\mathbf{x}_{i}),f_{\text{enc}}(\mathbf{d}))\leq 1-\epsilon\right\},$ (9)
@@ 4.2.3 Adversarial Contextual Fine-Tuning
$\mathcal{L}_{\text{adv}}=\frac{1}{N}\sum_{i=1}^{N}\text{CE}\left(F_{\theta_{T}}(\mathbf{x}_{i},\mathcal{D}_{\mathbf{x}_{i}}^{\Delta}),y_{i}\right),$ (10)
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{train}}+\lambda\cdot\mathcal{L}_{\text{adv}},$ (11)
@@ 4.3 Selective Knowledge Distillation with Evolving Knowledge Base
Low-latency intent detection, a crucial requirement for real-time applications, motivates the use of compact models; however, as shown in Appendix A.2, compact models often struggle to capture context-dependent patterns and generalize to unseen domains. Knowledge distillation partially mitigates this gap by transferring knowledge from a well-trained larger teacher model to a compact student model; however, in context-augmented settings, this method often fails to preserve contextual reasoning, and student models may degrade without continuously updated knowledge. To address these challenges, ADRAG introduces a unified deployment framework combining Selective Knowledge Distillation (SKD) with an Evolving Knowledge Base (EKB). SKD enables controllable and adaptive knowledge transfer through a encoder scheduler that gradually shifts supervision from teacher to student, enhancing transfer effectiveness. Meanwhile, EKB supplies real-time, query-specific context to support continual adaptation. Together, these components ensure accurate and efficient malicious intent detection in rapidly changing environments.
@@ 4.3.1 Selective Knowledge Distillation
Traditional knowledge distillation faces a fundamental tradeoff between stability and adaptability. Offline distillation relies on fixed teacher outputs, providing stable knowledge transfer but often failing to capture subtle, context-dependent, or adversarial intent. In contrast, online distillation introduces dynamic between teacher and student, improving adaptability but reducing controllability and destabilizing training, especially when the teacher is under-trained or over-fitted. To address this limitation, ADRAG introduces a encoder scheduler, a dynamic mechanism that governs the training process by assigning supervision modes to each epoch, defined as
$\mathbf{S}_{\text{train}}=[s_{1},s_{2},\dots,s_{T}],\quad s_{t}\in\{0,1,2\},$ (12)
@@
In the first stage with $s_{t}=0$, the scheduler activates only the teacher, optimizing it with the loss in Eq. 11. The student remains inactive to avoid premature exposure to unstable supervision signals, reducing noise propagation and establishing a stable foundation for knowledge transfer.
@@
$\mathcal{L}_{\text{S}}^{(b)}=\mathrm{KL}\left(F_{\theta_{T}^{(b)}}(\cdot\mid\mathbf{x}_{b},\mathcal{D}_{\mathbf{x}_{b}})\,\|\,F_{\theta_{S}^{(b)}}(\cdot\mid\mathbf{x}_{b},\mathcal{D}_{\mathbf{x}_{b}})\right),$ (13)
@@
$\mathcal{L}_{\text{S}}=\mathrm{KL}\left(F_{\theta_{T}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\,\|\,F_{\theta_{S}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\right).$ (14)
@@ 4.3.2 Evolving Knowledge Base
$\text{Conf}(\{y_{i}^{(j)}\})=\mathbb{I}\left[\forall j,\,y_{i}^{(j)}=y_{i}\right]\cdot\mathbb{I}[n\geq k],$ (15)
@@
$(x_{i},y_{i})\sim F_{\theta_{SG}}\left(p(y_{i}),\{e_{1},e_{2},\dots,e_{k}\}\right),$ (16)
@@ 5 Experiments
We evaluate ADRAG against three representative approaches for malicious intent detection: (1) a supervised fine-tuned BERT classifier; (2) GPT-4 (OpenAI et al., 2023), evaluated with three methods and diverse prompting (details provided in Appendix A.8); and (3) two dedicated safety LLMs, including Llama-Guard-3 (Inan et al., 2023) from Meta and WildGuard (Han et al., 2024) from AllenAI.
@@
The ADRAG base models are ModernBERT-395M (teacher) and ModernBERT-149M (student)  (Warner et al., 2024). Models are fine-tuned with PyTorch (Paszke et al., 2019), Hugging Face Transformer (Wolf et al., 2020) and DeepSpeed (Rasley et al., 2020) on NVIDIA 70 H200 nodes. Retrieval uses FAISS, prepending top-$K$ exemplars to each query. Adversarial Knowledge Base Generation and Policy-Guided Synthetic Generation leverage the uncensored Dolphin-2.9.1-Mistral-22B (theo77186, 2024). Hyperparameters (batch size, learning rate, retrieval $K$, data weights, distillation scheduler) are tuned via Weights & Biases on validation data. Deployment uses Ray (Moritz et al., 2018) and vLLM (Kwon et al., 2023) for high-throughput, low-latency serving. Performance is reported using weighted F1 score for classification accuracy and queries per second (QPS) for latency.

### v2

@@ 1 Introduction
Beyond these limitations, most traditional methods are designed for offline use and cannot effectively support real-time, context-aware intent detection due to architectural and latency constraints. For example, a music application query such as “songs like ‘I Deserve to Bleed’", intended for search and recommendation, may be incorrectly flagged as unsafe if the model lacks knowledge that “I Deserve to Bleed” is a song title. In practice, enabling such capabilities requires overcoming two main challenges: (1) lightweight models may meet latency requirements but often lack the reasoning and generalization needed for nuanced or adversarial queries; and (2) effective moderation increasingly requires dynamic, input-specific context that static classifiers cannot handle.
In this paper, we present ADRAG (Adversarial Distilled Retrieval-Augmented Guard), a unified framework for accurate, robust, and efficient online malicious intent detection. In the training stage, a high-capacity teacher model is trained on adversarially perturbed, retrieval-augmented inputs to learn context-aware classification under adversarial conditions. In the inference stage, a distillation encoder scheduler transfers the teacher’s knowledge to a compact student guard model, whose capabilities are continually refined through an online knowledge base. At deployment, the student model retrieves top-$K$ safety examples from the knowledge base to enable real-time, context-aware malicious intent detection.
@@
    We demonstrate that ADRAG, with a 149M-parameter model, matches the performance of GPT-4, WildGuard-7B, and Llama-Guard-3-8B across ten safety benchmarks, while sustaining an online latency under 6 ms at 300 QPS.
@@ 2 Related Work
Retrieval-augmented generation (RAG) enhances grounding and robustness by incorporating external context (Lewis et al., 2020). Beyond generation, recent work has extended RAG for safety tasks. Class-RAG (Chen et al., 2024) fine-tune Llama-3-8b with retrieved examples to adapt to changing moderation policies, while GuardAgent (Xiang et al., 2024) leverages retrieval to enforce safety constraints in planning agents. However, retrieved content may introduce harmful or misleading information (An et al., 2025). While RAG improves adaptability, its retrieval component introduces new latency challenges. To address this, Knowledge distillation (Hinton et al., 2015) trains compact student models to approximate the behavior of larger teacher models. This approach has enabled efficient NLP models such as DistilBERT (Sanh et al., 2019), and MiniLM (Wang et al., 2020). In the AI safety domain, distillation has been applied to build lightweight systems such as HarmAug (Zhang et al., 2024c), which distills Llama-3-8B into a 435M-parameter model. Our work combines RAG’s adaptability with distillation’s efficiency to deliver accurate, robust, and controllable malicious intent detection for real-time deployment.
@@ 3.1 Large Language Models
$F_{\theta}(\mathbf{y}\mid\mathbf{x})=\prod_{t=n+1}^{T}p(x_{t}\mid x_{<t},\mathbf{x}),$ (3.1)
@@ 3.2 Retrieval-Augmented Generation
$\mathcal{D}_{\mathbf{x}}=\mathcal{R}_{K}(\mathbf{x};\mathcal{D}),$ (3.2)
@@ 3.3 Adversarial Training
$\min_{\theta}\;\mathbb{E}_{\mathbf{x},\mathbf{y}}\left[\max_{\|\delta\|\leq\rho}-\log F_{\theta}^{\delta}(\mathbf{y}\mid\mathbf{x})\right],$ (3.3)
@@ 3.4 Knowledge Distillation
$\mathcal{L}_{\text{distill}}=\mathrm{KL}\left(F_{\theta_{T}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\,\|\,F_{\theta_{S}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\right),$ (3.4)
@@ 3.5 Problem Statement
$\hat{y}=F_{\theta}(\mathbf{x},\mathcal{D}_{\mathbf{x}}),\quad\mathcal{T}_{\text{tot}}(\mathbf{x})=\mathcal{T}_{\text{ret}}(\mathbf{x})+\mathcal{T}_{\text{inf}}(\mathbf{x},\mathcal{D}_{\mathbf{x}})\leq\tau,$ (3.5)
where $\mathcal{T}_{\text{tot}}$ denotes the total system latency, comprising retrieval latency $\mathcal{T}_{\text{ret}}$ and inference latency $\mathcal{T}_{\text{inf}}$, which need to remain below a strict threshold $\tau$ to enable real-time moderation. To support context-aware prediction, the query $\mathbf{x}$ is concatenated with $K$ retrieved examples: $\mathbf{x};\mathbf{d}_{\mathbf{x}}^{(1)};\ldots;\mathbf{d}_{\mathbf{x}}^{(K)}]$, where each $\mathbf{d}_{\mathbf{x}}^{(i)}=(\bar{\mathbf{x}}_{i},\bar{y}_{i})$ represents a prior query and its associated intent label. The concatenated input is fed into the classifier $F_{\theta}$ (or $\mathbf{x}$ alone if $\mathcal{D}_{\mathbf{x}}=\emptyset$), with prompt engineering used to ensure a coherent sequence that preserves semantic clarity and guides model behavior. This context-enhanced formulation improves reliability by leveraging semantically similar examples, while maintaining the strict latency constraints required for real-time applications.
@@ 4.1 Overview
We propose ADRAG, a two-stage framework for Accurate, Robust, and Efficient online malicious intent detection, as illustrated in Figure 1. In the training phase, a high-capacity teacher model $F_{\theta_{T}}(\mathbf{x},\mathcal{D}_{\mathbf{x}})$ is optimized via Retrieval-Augmented Adversarial Fine-Tuning (RAFT), where each input $\mathbf{x}$ is paired with a retrieved context set $\mathcal{D}_{\mathbf{x}}$ that injects external, often domain-specific knowledge to enrich under-specified queries, substantially enhancing the model’s predictive accuracy, especially for ambiguous or information-sparse cases. To further enhance robustness, ADRAG simulates real-world noise in the knowledge base, retrieval pipeline, context quality, and user input through adversarial perturbations, incorporating them into training via an adversarial loss $\mathcal{L}_{\text{adv}}$ that promotes consistent and reliable predictions under degraded conditions. In the inference phase, ADRAG applies Selective Knowledge Distillation (SKD) to compress the teacher into a lightweight student model $F_{\theta_{S}}$, guided by an encoder scheduler that governs training dynamics by assigning each epoch to operate in teacher-only, co-distillation, or student-focused mode. At deployment, the student retrieves top-$K$ relevant contexts $\mathcal{D}_{\mathbf{x}}\subset\mathcal{D}$ for each query $\mathbf{x}$ to enable accurate, low-latency predictions that generalize beyond the static training distribution and remain resilient to adversarial or ambiguous inputs, while adaptability is maintained through an Evolving Knowledge Base (EKB) continuously updated with user feedback and policy-driven synthetic examples.
@@ 4.2.1 Supervised Contextual Fine-Tuning
$\mathcal{D}_{\mathbf{x}_{i}}=\left\{\mathbf{d}\in\mathcal{R}_{K}(\mathbf{x}_{i};\mathcal{D})\;\middle|\;\text{sim}(f_{\text{enc}}(\mathbf{x}_{i}),f_{\text{enc}}(\mathbf{d}))\geq 1-\epsilon\right\},$ (4.1)
@@
$\mathcal{L}_{\text{train}}=\frac{1}{N}\sum_{i=1}^{N}\text{CE}\left(F_{\theta_{T}}(\mathbf{x}_{i},\mathcal{D}_{\mathbf{x}_{i}}),y_{i}\right),$ (4.2)
@@
$\mathcal{D}_{\mathbf{x}_{i}}^{\Delta}=\mathcal{R}_{K}(\mathbf{x}_{i};\mathcal{D},f_{\text{enc}}^{\Delta},\text{sim}^{\Delta}),$ (4.3)
@@
$\mathcal{D}_{\mathbf{x}_{i}}^{\Delta}=\left\{\mathbf{d}\in\mathcal{D}\;\middle|\;\delta\leq\text{sim}(f_{\text{enc}}(\mathbf{x}_{i}),f_{\text{enc}}(\mathbf{d}))\leq 1-\epsilon\right\},$ (4.4)
@@ 4.2.3 Adversarial Contextual Fine-Tuning
$\mathcal{L}_{\text{adv}}=\frac{1}{N}\sum_{i=1}^{N}\text{CE}\left(F_{\theta_{T}}(\mathbf{x}_{i},\mathcal{D}_{\mathbf{x}_{i}}^{\Delta}),y_{i}\right),$ (4.5)
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{train}}+\lambda\cdot\mathcal{L}_{\text{adv}},$ (4.6)
@@ 4.3 Selective Knowledge Distillation with Evolving Knowledge Base
Low-latency intent detection, a crucial requirement for real-time applications, motivates the use of compact models; however, as shown in Appendix A.2, compact models often struggle to capture context-dependent patterns and generalize to unseen domains. Knowledge distillation partially mitigates this gap by transferring knowledge from a well-trained larger teacher model to a compact student model; however, in context-augmented settings, this method often fails to preserve contextual reasoning, and student models may degrade without continuously updated knowledge. To address these challenges, ADRAG introduces a unified deployment framework combining Selective Knowledge Distillation (SKD) with an Evolving Knowledge Base (EKB). SKD enables controllable and adaptive knowledge transfer through an encoder scheduler that gradually shifts supervision from teacher to student, enhancing transfer effectiveness. Meanwhile, EKB supplies real-time, query-specific context to support continual adaptation. Together, these components ensure accurate and efficient malicious intent detection in rapidly changing environments.
@@ 4.3.1 Selective Knowledge Distillation
Traditional knowledge distillation faces a fundamental tradeoff between stability and adaptability. Offline distillation relies on fixed teacher outputs, providing stable knowledge transfer but often failing to capture subtle, context-dependent, or adversarial intent. In contrast, online distillation introduces dynamic between teacher and student, improving adaptability but reducing controllability and destabilizing training, especially when the teacher is under-trained or over-fitted. To address this limitation, ADRAG introduces an encoder scheduler, a dynamic mechanism that governs the training process by assigning supervision modes to each epoch, defined as
$\mathbf{S}_{\text{train}}=[s_{1},s_{2},\dots,s_{T}],\quad s_{t}\in\{0,1,2\},$ (4.7)
@@
In the first stage with $s_{t}=0$, the scheduler activates only the teacher, optimizing it with the loss in Eq. 4.6. The student remains inactive to avoid premature exposure to unstable supervision signals, reducing noise propagation and establishing a stable foundation for knowledge transfer.
@@
$\mathcal{L}_{\text{S}}^{(b)}=\mathrm{KL}\left(F_{\theta_{T}^{(b)}}(\cdot\mid\mathbf{x}_{b},\mathcal{D}_{\mathbf{x}_{b}})\,\|\,F_{\theta_{S}^{(b)}}(\cdot\mid\mathbf{x}_{b},\mathcal{D}_{\mathbf{x}_{b}})\right),$ (4.8)
@@
$\mathcal{L}_{\text{S}}=\mathrm{KL}\left(F_{\theta_{T}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\,\|\,F_{\theta_{S}}(\cdot\mid\mathbf{x},\mathcal{D}_{\mathbf{x}})\right).$ (4.9)
@@ 4.3.2 Evolving Knowledge Base
$\text{Conf}(\{y_{i}^{(j)}\})=\mathbb{I}\left[\forall j,\,y_{i}^{(j)}=y_{i}\right]\cdot\mathbb{I}[n\geq k],$ (4.10)
@@
$(x_{i},y_{i})\sim F_{\theta_{SG}}\left(p(y_{i}),\{e_{1},e_{2},\dots,e_{k}\}\right),$ (4.11)
@@ 5 Experiments
We evaluate ADRAG against three representative approaches for malicious intent detection: (1) A supervised fine-tuned ModernBERT (Warner et al., 2024), optimized through 600 hyperparameter trials; (2) GPT-4 (OpenAI et al., 2023), evaluated with three methods and diverse prompting; and (3) Dedicated safety LLMs, including Llama-Guard-3 (Inan et al., 2023) from Meta and WildGuard (Han et al., 2024) from AllenAI. Details of the baselines setup and evaluation are provided in Appendices A.8, A.10, A.11, and A.12.
@@
The ADRAG base models are ModernBERT-395M (teacher) and ModernBERT-149M (student) (Warner et al., 2024). Models are fine-tuned with PyTorch (Paszke et al., 2019), Hugging Face Transformer (Wolf et al., 2020) and DeepSpeed (Rasley et al., 2020) on 70 NVIDIA H200 nodes. Retrieval uses FAISS, prepending top-$K$ exemplars to each query. Adversarial Knowledge Base Generation and Policy-Guided Synthetic Generation leverage the uncensored Dolphin-2.9.1-Mistral-22B (theo77186, 2024). Hyperparameters (batch size, learning rate, retrieval $K$, data weights, distillation scheduler) are tuned via Weights & Biases on validation data. Deployment uses Ray (Moritz et al., 2018) and vLLM (Kwon et al., 2023) for high-throughput, low-latency serving. Performance is reported using weighted F1 score for classification accuracy and queries per second (QPS) for latency.